A dangerous pollutant present in Laguna de Bay is mercury (II) chloride. Scientists are suggesting that

the local government use aluminum to remove the dangerous compound and trap the free mercury

formed. This is an example of a single replacement reaction involving Al +3 , Cl -1 and Hg +2 . Give the

balanced chemical equation for the reaction. If you have approximately 200 milligrams of mercury (II)

chloride in a sample container and used only 10 milligrams of the aluminum ion, how many mercury

ions are you going to produce and how much of the pollutant will be left?


1
Expert's answer
2021-02-18T07:45:09-0500

2Al + 3HgCl2 ---> 3Hg + 2AlCl3

n(HgCl2) = 200/271,52/3 = 0.245 moles

n(AlCl3) = 10/133/2=0.0376 moles.

n(HgCl2) - n(AlCl3) = 0.207 moles.

Solution: 0.207 moles of HgCl2 will be left.